Jamie and her friends brought 1 cheese pizza, 1 mashroom pizza and 1 pepperoni pizza. The cost of each include tax are:
Cheese $7.95
Mashroom $8.95
Pepporni 8.45
Each Pizza consit 0f 8 slices. Jamie ate 1 slice of cheese, 2 slices of mashroom, and 1 slice of pepporni. What was the cost of the pizza that Jamie ate to the nearest tenth?

Answer by tutorcecilia(2152)

Find the cost per slice:
Cheese $7.95/8 = $0.9375 per slice
Mashroom $8.95/8 = $1.11875 per slice
Pepporni 8.45/8 = $1.05625 per slice
Jamie ate:
1 slice of cheese x $0.9375 = $0.9375
2 slices of mashroom x $1.11875 = $2.2375
1 slice of pepporni x $1.05625
___________________________________________
TOTAL AMOUNT OF PIZZA JAMIE ATE: $4.23125 = $4.23